# Anisotropy can make a moving active fluid membrane rough or crumpled

### v1

@@ Abstract
Figure 4: (a) Different phases and the RG flow lines of $\alpha$ in $\beta$-$\alpha$ plane. Yellow regions are the logarithmic rough phase. Central white region is the KPZ like putative algebraic rough phase. The black lines in the rough phase show the locus of the fixed point value $\alpha^{*}$ as a function of $\beta$. (b) The RG flow diagram in the $\beta$-$\alpha$-$g$ space. Arrows pointing toward (away from) the $\beta$-$\alpha$ plane represent the direction of RG flows, indicating stable (unstable) regions. Green regions in (a) and (b) correspond to crumpled phase (see text).
Figure 4(a) shows the phases and the RG flow lines of $\alpha$ in $\beta$-$\alpha$ plane. Since $d\beta/dl=0$, the RG flow lines in the $\beta$-$\alpha$ plane are perpendicular to the $\beta$-axis. The green horizontal regions at the top and bottom of Fig. 4(a) have $|\alpha|>1$, corresponding to the crumpled phase. The remaining region of this figure with $|\alpha|<1$, corresponding to linear stability, has two parts: The regions with $\beta^{2}>1$ (marked yellow) are perturbatively accessible logarithmically rough phase, whereas the region with $\beta^{2}<1$ is perturbatively inaccessible. Figure 4(b) shows the RG flow lines in the $\beta$-$\alpha$-$g$ space. Flow lines with initial values of $g$ in the yellow region of $\alpha$-$\beta$ plane flow to zero giving logarithmic roughness. Flow lines with initial values of $g$ inside the whitish cuboid region flow out to the perturbatively inaccessible strong coupling fixed point indicating KPZ-like algebraic roughness.
In our DNS studies, used to corroborate our RG results and characterize the phases, we discretize Eq. (11) by standard forward-backward differences on a square grid with lattice spacing $a_{0}=1$, and numerically solve it by using the Euler algorithm; see SM sup for more details. We calculate ${\mathcal{W}}$ for model parameters falling in the (i) yellow, stable region (logarithmically rough), (ii) unstable, white region (speculated to be algebraically rough) and (iii) green, linearly unstable (crumpled) region in the phase diagram 4(a). For the logarithmic rough phase, we have set $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=2$ and $\overline{\lambda}=4$ such that $|\alpha|<1$ and $|\beta|>1$ and system sizes $L=100$, $300$, $500$, $800$, $1000$ and $1600$ and time step $\Delta t=0.05$ in our simulations. Numerical measurements of ${\cal W}_{\text{sat}}$ for various $L$ give $\tilde{\chi}\approx 0.43\pm 0.019$, close to the analytically obtained value $\tilde{\chi}=0.50$, firmly establishing the logarithmic rough phase.
@@
Next we set $\nu=0.50$, $\overline{\nu}=0.25$, $\lambda=24$ and $\overline{\lambda}=1$ such that $|\alpha|<1$ and $|\beta|<1$, corresponding to $dg(l)/dl>0$ [see Eq. (9)]. Time step $\Delta t=0.01$. With ${\cal W}\sim t^{\tilde{\beta}}$ initially, we obtained $\tilde{\beta}\approx 0.234\pm 1.644\times 10^{-5}$; see also SM sup . Furthermore, numerically measuring ${\mathcal{W}}_{\text{sat}}$ for $L=100$, $200$, $300$, $400$ and $500$ give $\chi\approx 0.379\pm 0.008$, and hence $z=\chi/\tilde{\beta}=1.619$ with $\chi+z\approx 2$ (as it should be). Our results on $\chi$ confirms an algebraically rough phase with translational SRO and orientational LRO, as speculated from our RG calculations above. The value of $\chi=0.379$ shows reasonable agreement with the existing results Oliveira et al. (2013); Aarão Reis (2004); Kondev et al. (2000); Pagnani and Parisi (2015); Kelling et al. (2017); Ódor et al. (2009); Halpin-Healy (2012); Chin and den Nijs (1999); Colaiori and Moore (2001); Gomes-Filho et al. (2021) and also with a mode coupling theory prediction Bhattacharjee (1998), indicating scaling behavior statistically same as in the strong coupling rough phase of the 2D KPZ equation. These DNS results lend further credence to a strong coupling rough phase with large-scale isotropy.
@@
So far we have considered a nearly flat patch within a Monge parametrization, which holds up to length scales at which curvature $C$ becomes important. Within this scale, while microscopically the membrane is inversion-asymmetric, the effective long wavelength behavior is symmetric in the logarithmic rough phase. The nonlinear terms with coefficients $\lambda_{ij}$ break the inversion symmetry. However, coupling $g$ vanishes in the long wavelength limit (albeit slowly enough to introduce logarithmic modulations to the noise and damping). At larger scales, when $C$ should be important, the spontaneous curvature term scales as $C/r$, whereas the surface tension term scales as $r^{0}$, making $C$ increasingly unimportant (relative to the surface tension) at larger length scales. We thus expect the membrane to remain effectively inversion-symmetric (and statistically flat) at arbitrarily large scales. In the algebraic rough phase, $g$ is not expected to vanish, giving inversion-asymmetry to survive in the long wavelength limit. Nonetheless, with expected $z<2$ the surface tension continues to dominate over $C$, ultimately giving a statistically flat but inversion-asymmetric membrane. At the onset of the crumpled phase, the effective surface tension vanishes, making $C$ important. This is beyond the scope of the present study.
@@
We have thus shown that the hydrodynamic theory for a moving active anisotropic inversion-asymmetric fluid membrane is given by an anisotropic KPZ equation, giving a logarithmically rough phase an algebraically rough phase and a crumpled phase. The average membrane velocity $\lambda_{ij}\langle{\nabla_{i}h}\nabla_{j}h\rangle$, an observable feature and proportional to $g$, vanishes in the inversion-symmetric logarithmic rough phase, but is finite in the asymmetric algebraically rough phase. It would be interesting to explore connections between our predicted crumpled phase and observed reduction of membrane tension to even negative values during membrane fusion Solon et al. (2006). Membrane tension may be measured using optical set ups Popescu et al. (2006), giving information about $\nu,\overline{\nu}$. Comparison with experimental results on membrane fluctuations by using, e.g., spectroscopy Betz and Sykes (2012) can yield information about $\lambda,\overline{\lambda}$. Studying how the fluid flows along the membrane in the different phases and their dependence on the tension anisotropy are related to the cortical flows reported in Ref. Mayer et al. (2010) should be promising future work directions.
For reasons of simplicity, we set $\nu_{xx}=\nu_{yy}$ and $\lambda_{xx}=\lambda_{yy}$. We note that $\nu_{xx}\neq\nu_{yy}$ together with $\lambda_{xx}\neq\lambda_{yy}$ (but without $\overline{\nu},\overline{\lambda}$) was considered in Refs. Wolf (1991); Altman et al. (2015), giving QLRO translational order and 2D KPZ-like instability for different choices of an effective anisotropy parameter. We thus speculate that even in the general case with $\nu_{xx}\neq\nu_{yy}$, $\lambda_{xx}\neq\lambda_{yy}$ and with non-vanishing $\overline{\nu},\overline{\lambda}$ can still have just the three phases studied here.It will be interesting to connect Eq. (11) to variants of the Toom model Barabási et al. (1992); Jeong et al. (1993), and generalize our study by considering the interplay between the nonlocal effects Jana et al. (2024); Jana and Basu (2024) and anisotropy.
@@
Appendix A Linear instability analysis
We start from Eq. (1) of the main text for an anisotropic membrane with a fixed background, ignoring the conservation of embedding bulk fluid momentum. As argued, the hydrodynamic equation governing the undulation modes of such a membrane, described by the height field $h({\bf x},t)$, is an anisotropic Kardar-Parisi-Zhang (KPZ) equation with a Gaussian-distributed, zero-mean white noise $f({\bf x},t)$
$\frac{\partial h}{\partial t}+\frac{\lambda_{ij}}{2}\partial_{i}h\partial_{j}h=\nu_{ij}\partial_{i}\partial_{j}h+f.$ (11)
Here, both $i$ and $j$ can take value $1$ or $2$. To perform linear stability analysis, we consider only the linear terms in Eq. (11), i.e., set $\lambda_{ij}=0$. The resulting linearized equation of $h$ in Fourier space is
$\partial_{t}h=-\nu(k_{x}^{2}+k_{y}^{2}+2\alpha k_{x}k_{y})h,$ (12)
[…]
When $(1+\alpha\sin 2\theta)<0$, linear instability ensues. The growth rate of initial fluctuations is maximized when $(1+\alpha\sin 2\theta)$ attains its largest negative value. For $\alpha>1$, this happens at $\theta=3\pi/4$, corresponding to $k_{x}=-k_{y}$ giving the Fourier wavevector with the maximum growth rate. Keeping only the corresponding Fourier mode (which is the most dominant Fourier mode), we get

### v2

@@ Abstract
Figure 2: (a) Two distinct phases in $\beta$-$\alpha$ plane. White region is the KPZ like putative algebraic rough phase. (b) The RG flow diagram in the $\beta$-$\alpha$-$g$ space. Arrows pointing away from the $\beta$-$\alpha$ plane represent the direction of RG flows, indicating a strong coupling algebraic rough phase. Green regions in (a) and (b) correspond to crumpled phase (see text).
We now theoretically explain the above numerical observations, quantitatively characterizing the phases, by using Eq. (1). Equation (1) admits a pseudo-Galilean invariance under the transformation $t^{\prime}=t,\,x_{i}^{\prime}=x_{i}+\lambda u_{i}t,\,h^{\prime}({\bf x}^{\prime},t^{\prime})=h({\bf x},t)+u_{i}x_{i}+\frac{\lambda}{2}u^{2}\,t$ sup . In (1), $\nu_{ij}$ plays the role of an (active) anisotropic surface tension. With $\nu_{x},\nu_{y}>0$, linear stability requires that $\nu_{x}\nu_{y}\geq\overline{\nu}^{2}$, i.e. $\beta\geq\alpha^{2}$; otherwise instability ensues. Here $\beta=\frac{\nu_{y}}{\nu_{x}}$ and $\alpha=\frac{\overline{\nu}}{\nu_{x}}$. If $\alpha=0$, Eq. (1) is invariant under (separately) $x\rightarrow-x$ or $y\rightarrow-y$; for $\alpha\neq 0$, the invariance is only under the joint inversion of $x$ and $y$. This means $\alpha_{c}=\pm\sqrt{\beta}$ is a threshold of $\alpha$, with larger $|\alpha|$ implies instability. This is the anisotropy-induced instability mentioned earlier in this Letter, which can occur even at zero noises, provided $|\alpha|>\sqrt{\beta}$. This then means in general the variance of the local height fluctuations $\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle$ and local normal fluctuation $\delta{\bf n}=-{\bm{\nabla}}h$, $\langle(\delta{\bf n}({\bf x},t))^{2}\rangle$ diverge. Can this instability be suppressed by including stabilizing higher order damping terms, e.g., $\nu_{4}\nabla^{4}h,\,\nu_{4}>0$ in (1)? Such a term can be generically present, but is irrelevant if there is no linear instability. In the event of linear instability, these higher order (in gradient) stabilizing terms are important. Their presence would ensure that at high enough wavevectors the system is linearly stable: If $\nu_{1},\nu_{2}$ are the eigenvalues of of $\nu_{ij}$, then for membranes of linear size $L<(>)L_{c}\sim\sqrt{\nu_{4}/|\text{min}(\nu_{1},\nu_{2})|}$ are dominated by the stabilizing hyperdiffusion (destabilizing diffusion) terms. In this case, $\langle[h({\bf x},t)-\overline{h}(t)]^{2}\rangle$ and $\delta{\bf n}=-{\bm{\nabla}}h$, $\langle(\delta{\bf n}({\bf x},t))^{2}\rangle$ diverge not at every scale, but only if the linear system size $L$ exceeds a threshold that scales with $\sqrt{\nu_{4}}$ Mukherjee and Basu (2024). Such divergences of $\langle(\delta{\bf n}({\bf x},t))^{2}\rangle$ for sufficiently larger membranes indicates that such membranes with $L>L_{c}$ exceeding the instability threshold may be crumpled, although (1) cannot be used to study the crumpled phase; see also later and Ref. Mukherjee and Basu (2022). This instability is reminiscent of the strongly crumpled phase in tensionless equilibrium asymmetric tethered membranes Banerjee et al. (2019a, b). In this unstable phase, the variance of the local normal fluctuations $\sim\langle({\bm{\nabla}}h)^{2}\rangle$ diverges for a sufficiently large but finite $L$, giving translational and orientational SRO. In this phase, the growth rate is maximum forming ripples at a particular angle depending on parameters $\nu_{x}$, $\nu_{y}$ and $\overline{\nu}$; see Fig. 1(c). Consider the special case where $\nu_{x}=\nu_{y}$ and $|\alpha|>1$ then ripples will form at an angle $\pi/4$ ($3\pi/4$) with the $x$-axis for $\alpha>1$ ($\alpha<-1$); see SM sup . See the ripples in Fig. 1(d), forming at an angle $\pi/4$, and the corresponding movies MOV2 and MOV3 in SM sup .
We now turn to the stable case with $|\alpha|<\sqrt{\beta}$, and consider its stability and fluctuation properties. In the linear theory with $|\alpha|<\sqrt{\beta}$, width ${\mathcal{W}}_{\text{sat}}\sim\sqrt{\ln(L/a_{0})}$. As in the usual KPZ equation Kardar et al. (1986), the combination of noises and nonlinearities here can change the linear theory results substantially. To study the nonlinear effects in the present model (1), we perform a one-loop perturbative dynamic renormalization group (RG) analysis of Eq. (1) with $|\alpha|<\sqrt{\beta}$. The Wilson dynamic RG method Täuber (2014); Forster et al. (1977) for our model closely resembles that for the KPZ equation Kardar et al. (1986); Barabási and Stanley (1995); Täuber (2014). Galilean invariance of (1) ensures that $\lambda$ is unrenormalized, which is an exact statement and the analogue of the nonrenormalization of the nonlinear coupling in the usual KPZ equation Barabási and Stanley (1995); Frey and Täuber (1994); see also SM sup . However, there are diverging one-loop corrections to $\nu_{x},\,\,\nu_{y},\,\overline{\nu}$ and $D$, which are represented by the one-loop Feynman diagrams (see SM) sup . Dimensional analysis allows us to define an effective dimensionless coupling constant $g$ having a critical dimension 2 as
$g=\frac{\lambda^{2}D}{16\pi(\nu_{x}\nu_{y}-\overline{\nu}^{2})^{\frac{3}{2}}},$ (6)
along with $\alpha$, $\beta$ defined earlier. The RG differential recursion relations ($\exp(l)$ is a dimensionless length) are
$\frac{dD}{dl}=D\bigg[z-2-2\chi+g\frac{3(1+\beta^{2})/4+\beta/2+\alpha^{2}}{(\beta-\alpha^{2})}\bigg],$ (7)
$\frac{d\nu_{x}}{dl}=\nu_{x}[z-2-g(1/2-\beta/2)],$ (8)
$\frac{d\nu_{y}}{dl}=\nu_{y}[z-2-g(1/2-1/(2\beta))],$ (9)
$\frac{d\overline{\nu}}{dl}=\overline{\nu}[z-2-g],$ (10)
$\frac{dg}{dl}=2g^{2},$ (11)
$\frac{d\alpha}{dl}=-g(\alpha+\alpha\beta)/2,$ (12)
$\frac{d\beta}{dl}=g(1-\beta^{2})/2.$ (13)
Here, $z$ and $\chi$ are the dynamic and roughness exponents, respectively Barabási and Stanley (1995). Equation (11) shows that $g(l)$ diverges as $l\rightarrow l_{c}\sim{\cal O}(1)$ from below, where $l_{c}$ is a nonuniversal model parameter-dependent length. Of course, we cannot follow all the way to $g(l)$ diverging; as soon as $g(l)$ becomes ${\cal O}(1)$, the perturbation theory breaks down and hints a strong coupling algebraic rough phase similar to the usual 2D KPZ equation Tang et al. (1990); Barabási and Stanley (1995). Assuming $g=g^{*}$ at the putative perturbatively inaccessible strong coupling fixed point (SCFP), the flow equations (12) and (13) for $\alpha$ and $\beta$ respectively yield $\beta=1$ and $\alpha=0$ as the stable fixed points. We therefore speculate that, in the long-wavelength limit ($l\to\infty$), $\nu_{x}\to\nu_{y}$ and $\overline{\nu}\to 0$, indicating that the system becomes effectively isotropic and belongs to the KPZ universality class.
Figure 3: Log-log plots of $\mathcal{W}$ versus $t$ (a) for $L=500$ with $\nu_{x}=0.50$, $\overline{\nu}=0.10$, $D=0.005$, $\lambda=24$, and different values of $\nu_{y}$. Inset: linear-scale plot for $L=300$. (b) For $L=500$ with $\nu_{y}=0.55$ and varying $\overline{\nu}$ (other parameters as in (a)). Inset: plot for $L=300$ in linear-scale. (c) For $\nu_{y}=0.55$, $\overline{\nu}=0.20$, (other parameters as in (a)) and system sizes $L=100$ to $500$. The red dashed line indicates a linear fit with slope $\tilde{\beta}=0.22\pm 2.6\times 10^{-5}$. (d) Log-log plot of Saturation width $\mathcal{W}_{\text{sat}}$ versus $L$ for the same parameters as in (c), with a linear fit yielding $\chi=0.38\pm 0.002$ (see text).
Figure 2(a) shows the phase diagram in the $\beta$-$\alpha$ plane. The green regions at the top and bottom, characterized by $|\alpha|>\sqrt{\beta}$, correspond to the crumpled phase driven by a linear instability. The intervening white region, defined by $|\alpha|<\sqrt{\beta}$, is linearly stable but perturbatively inaccessible within the RG framework. Figure 2(b) displays the corresponding RG flow lines in the full $\beta$-$\alpha$-$g$ space. Flow lines originating within the whitish paraboloid region (bounded by $|\alpha|<\sqrt{\beta}$) escape to a SCFP that is perturbatively inaccessible, reminiscent of the behavior in the 2D isotropic KPZ equation Barabási and Stanley (1995); Tang et al. (1990)
To complement the predictions of this RG analysis, we perform DNS of Eq. (1). We discretize the equation using standard forward-backward finite differences on a square grid with lattice spacing $a_{0}=1$, and integrate it numerically via the Euler algorithm with a fixed time step $\Delta t=0.01$; further implementation details are provided in the SM sup . We compute the interface width $\mathcal{W}$} for parameter sets lying in both the (i) speculated algebraically rough regime and (ii) linearly unstable (crumpled) regime of Fig. 2(a).
@@
Now we focus on the DNS results within the speculated algebraic rough regime, which is argued to be statistically identical to the isotropic 2D KPZ equation. We chose parameter space such that $dg(l)/dl>0$ [see Eq. (11)]. Fig. 3(a) shows the $\mathcal{W}(t)$ for various $\nu_{y}$ values with fixed $\nu_{x},\overline{\nu}$. The linear growth regime exhibits an approximately constant slope across all cases, suggesting that the growth exponent $\tilde{\beta}$ is insensitive to $\nu_{y}$. However, the $\mathcal{W}_{\text{sat}}$ increases with decreasing $\nu_{y}$, as lower $\nu_{y}$ (for a fixed $\nu_{x}$) drives the system closer to the instability threshold. See Fig. 2(a). Fig. 3(b) presents $\mathcal{W}(t)$ for different $\overline{\nu}$. Again, the growth exponent remains nearly unchanged, while $\mathcal{W}_{\text{sat}}$ increases with $|\overline{\nu}|$, consistent with approaching the unstable region at fixed $\nu_{x}$. See Fig. 2(a). Fig. 3(c) shows variation of $\cal W$} over time for different $L$. The red dashed line indicates the slope of the growth region obtained through linear fitting, yielding $\tilde{\beta}=0.22\pm 2.6\times 10^{-5}$. Furthermore, numerically measuring ${\mathcal{W}}_{\text{sat}}$ for $L=100$, $200$, $300$, $400$ and $500$ give $\chi\approx 0.38\pm 0.002$, and hence $z=\chi/\tilde{\beta}=1.72$ with $\chi+z\approx 2$ (as it should be). See Fig. 3(d). Our results on $\chi$ confirms an algebraically rough phase with translational SRO and orientational LRO, as speculated from our RG calculations above. The value of $\chi=0.38$ compares reasonably with the existing results Oliveira et al. (2013); Aarão Reis (2004); Kondev et al. (2000); Pagnani and Parisi (2015); Kelling et al. (2017); Ódor et al. (2009); Halpin-Healy (2012); Chin and den Nijs (1999); Colaiori and Moore (2001); Halpin-Healy (2013) and also not far from a mode coupling theory Bhattacharjee (1998) and functional RG Halpin-Healy (1990) predictions, indicating 2D KPZ-like scaling behavior lending further credence to a strong coupling rough phase belonging to the 2D KPZ universality class.
To explore the unstable region, we introduce a regularizing term $\nu_{4}\nabla^{4}h$ to Eq. (1). For DNS in the unstable region (i.e green region in Fig. 2(a), (b)) we choose $\nu_{4}=0.10$, $\Delta t=0.01$. However, the interface width still diverges rapidly after a short duration, possibly indicating that additional higher-order nonlinear terms, neglected in (1) by using scaling arguments, are now required to describe this region reliably. Fig. 1(c) [(d)], with $L_{c}=0.46$ $L_{c}=0.32$], both much smaller than the membrane size $L=200$, display the height snapshots just before divergence occurs.
@@
So far we have considered a nearly flat patch within a Monge parametrization, which holds up to length scales at which curvature $C$ becomes important. At larger scales, when $C$ could be important, the spontaneous curvature term scales as $C/r$. In the algebraic rough phase, $g$ should not vanish, with inversion-asymmetry surviving in the long wavelength limit. Nonetheless, with expected $z<2$, $\nu_{ij}$ continues to dominate over $C$ (in a scaling sense), ultimately giving a statistically flat but inversion-asymmetric membrane. For the crumpled phase, it is not possible to comment on the large-scale structure of the membrane.
We have thus shown that the hydrodynamic theory for a moving active anisotropic fluid membrane is given by an anisotropic KPZ equation, giving an algebraically rough phase, that in spite of the anisotropy belongs to the isotropic 2D KPZ universality class (and hence isotropy is en emergent symmetry of the model in the long wavelength limit) and an unstable phase that is speculated to be a crumpled phase. Existence of this putative crumpled phase is linked to sufficiently strong breaking of the inversion symmetry of the model. Our results show that the fluctuation-induced average membrane velocity $\frac{\lambda}{2}\langle(\nabla h)^{2}\rangle$, proportional to $g$ is finite in the algebraically rough phase. We are unable to comment on it in the unstable phase. In living cells, lamellipodia where actin filaments grow nearly normal to the membrane, realize our model via activity-driven permeation flows Cagnetta et al. (2018, ). Similarly, in vitro supported bilayers with actin nucleators exhibit ATP-driven fluctuations and forces Waechtler et al. (2024) are applicable to our model. Our results suggest that coarse-grained experiments, designed to explore long wavelength scaling properties, cannot detect microscopic anisotropy when the surface is stable.
@@
It would be interesting to explore connections between our predicted unstable phase and observed reduction of membrane tension to even negative values during membrane fusion Solon et al. (2006). Membrane tension may be measured using optical set ups Popescu et al. (2006), giving information about $\nu_{x},\nu_{y}$ and $\overline{\nu}$. Comparison with experiments on membrane fluctuations by using, e.g., spectroscopy Betz and Sykes (2012) can give $\lambda$. Studying how the fluid flows along the membrane in the different phases and their dependence on the tension anisotropy are related to the cortical flows reported in Ref. Mayer et al. (2010) and how the membrane dynamics may couple with a diffusing density living on the membrane Haldar et al. (2023a, b) should be promising future work directions. It will be interesting to connect Eq. (1) to variants of the Toom model Barabási et al. (1992); Jeong et al. (1993), and generalize our study by considering the interplay between the nonlocal effects Jana et al. (2024); Jana and Basu (2024) and anisotropy.
@@
Supplemental Material
I Linear instability analysis
We start from Eq. (1) of the main text for an anisotropic membrane with a fixed background, ignoring the conservation of embedding bulk fluid momentum. As argued in the main text, the hydrodynamic equation governing the undulation modes of such a membrane, described by the height field $h({\bf x},t)$, is an anisotropic Kardar-Parisi-Zhang (KPZ) equation with a Gaussian-distributed, zero-mean white noise $f({\bf x},t)$
$\frac{\partial h}{\partial t}+\frac{\lambda}{2}{({\bm{\nabla}}h)^{2}}=\nu_{ij}\partial_{i}\partial_{j}h+f.$ (14)
Here, both $i$ and $j$ can take value $1$ or $2$. To perform linear stability analysis, we consider only the linear terms in Eq. (1), i.e., set $\lambda=0$. The resulting linearized equation of $h$ in Fourier space is
$\partial_{t}h=-\nu_{x}(k_{x}^{2}+\beta k_{y}^{2}+2\alpha k_{x}k_{y})h,$ (15)
where ${\bf k}=(k_{x},k_{y})$ is a wavevector. We define $\beta\equiv\nu_{y}/\nu_{x}$, $\alpha\equiv{\overline{\nu}}/{\nu_{x}}$; see also the main text. To proceed further, we use the polar coordinates and introduce $k_{x}=k\cos\theta$ and $k_{y}=k\sin\theta$, where $k=|{\bf k}|$ and $\theta=\tan^{-1}(k_{y}/k_{x})$ is the polar angle. Now for linear stability, $\nu_{x}k^{2}(\cos^{2}\theta+\beta\sin^{2}\theta+\alpha\sin 2\theta)<0$ for all $\theta$. With $\nu_{x},\,\nu_{y}>0$; as assumed in the main text, we must have $(1+\beta\tan^{2}\theta+2\alpha\tan\theta)>0$ for all $\theta$ to ensure linear stability. This means $\beta\geq\alpha^{2}$. When $1+\beta\tan^{2}\theta+2\alpha\tan\theta<0$, linear instability occurs. The growth rate of the fluctuations is maximized when $1+\beta\tan^{2}\theta+2\alpha\tan\theta$ reaches its most negative value, which occurs at a specific $\theta$ determined by the values of $\alpha$ and $\beta$. Consequently, a stripe-like pattern emerges based on the value of $\theta$. A representative snapshot of this case is shown in Fig. 1(c) of the main text. Now we consider the special case $\nu_{x}=\nu_{y}=\nu$, i.e. $\beta=1$ and $\alpha=\overline{\nu}/\nu$. For linear stability, $\nu k^{2}(1+\alpha\sin 2\theta)<0$ for all $\theta$. With $\nu>0$, we must have $(1+\alpha\sin 2\theta)>0$ for all $\theta$ to ensure linear stability. Since the maximum of $|\sin\,2\theta|$ is unity, the previous condition in turn means $|\alpha|\leq 1$, or $\nu\pm\overline{\nu}>0$. When $(1+\alpha\sin 2\theta)<0$, linear instability again ensues. The growth rate of the fluctuations is maximized when $(1+\alpha\sin 2\theta)$ attains its largest negative value. For $\alpha>1$, this happens at $\theta=3\pi/4$, corresponding to $k_{x}=-k_{y}$ giving the Fourier wavevector with the maximum growth rate. Keeping only the corresponding Fourier mode (which is the most dominant Fourier mode), we get